Q1 The linear system Ax = b is given by:

x1−x2 + 4x3 = 7

4x1 + 2x2 –x3= 18,

x1 + 3x2+ x3 = 16,

has the solution x=(3, 4,2)T. Using the initial guess   x (0)=(1, 1,1)T

  1. Solve the above system as is using: Gauss-Seidel method. If the error increases, what does that mean and what should you do? (see b below)
  2. Condition the system so that convergence is secured and solve using the Gauss-Siedel method.

Q2: Find a system of linear algebraic equations Ax=b that needs pivoting and carry out the solution (using a limited number of significant digits, say 3)   as Without pivoting, With partial pivoting, With total pivoting and Make conclusions.
Answer #1

lij 2,-2, +413 4x, +2, - 13 - 15 2 + 38, + 1 = 16 2 (0) = (1,1,017 73) x (3,4,5) a by Can (1) by egh (2) (7+2 2 - 4 %) † (18-a diagonall dominants So we change the order of ean to get matrix. New arrangement in 42 + 2%, -13 = 18 x + 3x₂ + x3 = 16 Y,09 4 1 (3) 2(5 4 Iteration 3 119-1 X 3,1108 + 1917 ] 3.0938 1 116 - 310938 -1.9167 ] = 3.6682 7.lv (7-3,0938 +3.6032) - 11892